If there is $4.15 in the jar, and there are 4 dimes, how many quarters are in the jar?
Answer by jim_thompson5910(28595)

4 dimes are worth 4*0.1 = 0.4 dollars.

So there are 4.15 - 0.4 = 3.75 dollars worth of quarters

So there are 3.75/0.25 = 15 quarters

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------